# Unraveling the dynamical behaviors in a quasiperiodic mosaic lattice

### v1

@@ III Effective Hamiltonian in the large $\lambda$ limit
In the large quasiperiodic potential strength region, we have unveiled that the dynamics in the quasiperiodic mosaic lattice exhibits some unique properties, including differences in density distribution between even and odd sites, as well as time scales linearly depend on the strength of quasiperiodic potential. To understand these unique dynamical behaviors, we shall provide an explanation by using an effective Hamiltonian in the subsequent discussion. Suppose that the quasiperiodic potential is sufficiently large, surpassing the strength of the hopping terms. We can apply the Schrieffer-Wolf (S-W) transformation method to construct the effective Hamiltonian for the quasiperiodic mosaic model. The effective Hamiltonian can provide a clear physical insight for understanding the unique dynamical behaviors in the large quasiperiodic potential strength region.
@@ III.1 Derivation of the effective Hamiltonian
$\hat{H}=\textbf{c}^{\dagger}H\textbf{c},$
@@
The subscript of e or o indicates the Hamiltonian defined on the even or odd sites. The details about S-W transformation can be found in Appendix B. To the first order, $\hat{H}_{e}$ and $\hat{H}_{o}$ are decoupled and their dynamics seem to be independent of each other. Next, we will study the properties the effective Hamiltonian and show how to use it to understand the unique dynamical behaviors of quasiperiodic mosaic model.
@@ III.2 Properties of the effective Hamiltonian
First, we present the eigenstates and eigenvalues of the effective Hamiltonian, comparing them with the results obtained from exact diagonalization (ED). The results are illustrated in Fig. 5. In Fig. 5 (a), we depict the eigenvalues from the effective Hamiltonian and the quasiperiodic mosaic model, exhibiting a good agreement between them. It is evident that the eigenvalues of $\hat{H}_{o}$ contribute to the energy levels of ”band center” around $E=0$, while the eigenvalues of $\hat{H}_{e}$ contribute to the energy levels apart from the ”band center”. In Fig. 5 (b), we show the IPR for the effective Hamiltonian and quasiperiodic mosaic model. The effective Hamiltonian can capture the extended-localized properties of a large number of states. Although some eigenstates at the edge of the energy spectrum $\hat{H}_{o}$ deviate from the true eigenstates, we will show these eigenstates are localized and thus do not contribute to the dynamical behaviors.
Figure 5: The comparison of eigenvalues (a) and eigenstates (b) obtained from the effective Hamiltonian and the quasiperiodic mosaic model. The black points, blue points and red points are the results from $\hat{H}$, $\hat{H}_{o}$ and $\hat{H}_{e}$ respectively. We choose $\lambda=10$ and $L=377$. $n$ labels the $n_{th}$ eigenstate.
@@
Figure 6: (a) IPR for eigensatets of $\hat{H}_{o}$ with different system sizes. We choose $\lambda=10$ and $L=377,~{}610,~{}987$. (b) Finite-size scaling analysis for different eigenstates.
(ii) $\hat{H}_{o}$ distributes on the odd sites. Both the onsite potential and hopping term have been modulated by the quasiperiodic potential equally. As shown in Fig. 6 (a): the IPRs for the eigenstates of $\hat{H}_{o}$ in the middle of the energy spectrum is much smaller than that at the two sides of the spectrum, indicating that eigenstates in the middle and sides are extended and localized, respectively. To see it more clearly, in Fig. 6 (b), we perform finite-size scaling analysis on typical eigenstates in different regions using the moments $D_{2}$ (Eq.5) to characterize the distribution information of the eigenstate. For our analysis, we take the logarithm of both sides of the equation Eq.(5) and obtain
$-\ln(I_{2})/\ln(L)=-c/\ln(L)+D_{2}.$
We determine $D_{2}$ by conducting a linear fit to the curve in the two-dimensional space spanned by $-\frac{\ln(I_{2})}{\ln(L)}$ and $-\frac{1}{\ln(L)}$. The results show that typical eigenstates are extended in the middle and localized at two sides of the spectrum, indicating the existence of mobility edges in the spectrum of $\hat{H}_{o}$.
From the effective Hamiltonian, we gain additional insights: (i) All the extended eigenstates can be traced back to the term $\hat{H}_{o}$. Despite the fact that the energy width of the extended states narrows as $\lambda$ increases, the proportion of extended states remains constant with the quasiperiodic potential strength. (ii) Two types of localized eigenstates exist in the quasiperiodic mosaic model. They distribute on different sites and have different physical origins. The first type of localized eigenstates originates from $\hat{H}_{e}$. These states are distributed at sites where the onsite potential is nonzero. Considering an eigenstate $|\psi_{j}\rangle$ with a localized center at the $j$-th site, its eigenvalue can be approximated as $\lambda_{j}+\frac{2J^{2}}{\lambda_{j}}$ which is proportional to quasiperiodic potential strength $\lambda$ in the large quasiperiodic potential strength case. The second type of localized eigenstates originates from $\hat{H}_{o}$. These eigenstates mainly distribute on sites where the potential is zero. The eigenvalues of this eigenstates are inversely proportional to the quasiperiodic potential strength $\lambda$.
@@
III.3 Understand the wave packet dynamics from the effective Hamiltonian
@@
$e^{-i\hat{H}t}|\psi(0)\rangle$ $\approx e^{-i(\hat{H}_{e}+\hat{H}_{o})t}|\psi(0)\rangle$
@@
Since the wave packet dynamics is dominated by $\hat{H}_{o}$, the wave function only propagates along the odd sites in the dressed basis ($d^{\dagger}_{j}|0\rangle$). For the initial state with the Gaussian wave packet located at the center of lattice, the particle number distribution $n^{d}_{j}=\langle\psi(t)|\hat{d}^{\dagger}_{j}\hat{d}_{j}|\psi(t)\rangle$ out of the initial wave packet only distributes on the odd sites. Nevertheless, according to the S-W transformation Eq.(9), the creation (annihilation) operator on the even sites in the dressed basis can be related to the operators on both even and odd sites in the original basis. Up to the second order, the density distribution on the even sites in the original basis can be explicitly expressed as
@@
$\frac{J^{2}}{\lambda_{j}^{2}}(\langle\psi(t)|\hat{d}^{\dagger}_{j-1}\hat{d}_{j+1}|\psi(t)\rangle+\langle\psi(t)|\hat{d}^{\dagger}_{j+1}\hat{d}_{j-1}|\psi(t)\rangle),$
@@
From the expression of $\hat{H}_{o}$, we observe that $\frac{1}{\lambda}$ is an overall factor of the Hamiltonian of $\hat{H}_{o}$. It follows that the dynamics governed by $\hat{H}_{o}$ is invariant after a time scale $t\rightarrow t/\lambda$. This is consistent with the numerical results as shown in the inset of Fig.2. The wave packet dynamics governed by the effective Hamiltonian is also shown in Fig.1(e) with the cyan and magenta lines, exhibiting an excellent agreement with the wave packet dynamics of the original Hamiltonian. Our results demonstrate that the effective Hamiltonian can describe the dynamical behaviors well for the mosaic model in the large quasiperiodic potential strength region.
IV Summary
In summary, we investigated the dynamics of a quasiperiodic mosaic model and unravel its peculiar dynamical behaviors in the large quasiperiodic potential region, exhibiting obvious different density distribution on even and odd sites. Our results also demonstrate that time scales of dynamical evolution are determined by the strength of the quasiperiodic potential linearly. Applying the S-W transformation, we derived an effective Hamiltonian in the limit of large quasiperiodic potential strength. The effective Hamiltonians can be divided into two decoupled Hamiltonians $\hat{H}_{e}$ and $\hat{H}_{o}$ defined on the even and odd sites, respectively. While all eigenstates of $H_{e}$ are localized, the eigenstates of $\hat{H}_{o}$ include both localized and extended eigenstates separated by mobility edge. The effective model describes the dynamical behaviors well for the mosaic model in the large quasiperiodic potential region. Our study provides a clear picture for understanding the peculiar dynamical behaviors in the quasiperiodic mosaic lattice.
@@ Appendix A Two-point correlation function
and the correlation function can be written as
@@
Next we demonstrate how to calculate $\langle\hat{c}^{\dagger}_{a}\hat{c}_{b}\rangle$: the initial state $|\psi(0)\rangle$ are chosen as:
@@
In order to calculate the correlation function, we define two matrix $P^{A}$ and $P^{B}$. $P^{A}$ and $P^{B}$ are both $L\times(L/2+1)$ matrix and they satisfy:
@@ Appendix B S-W transformation
Then we can write down the Hamiltonian as
@@ Appendix C Effectiveness of unitary transformation
In this study, we investigate the impact of increasing the system size, denoted as $L$, on the effectiveness of the effective Hamiltonian under the same $\lambda$. Specifically, we choose $L=3\times 10^{3}$ and compare the eigenvalues and inverse participation ratios IPR obtained from the effective Hamiltonian and exact ED. The main results are depicted in Fig.7.
@@
Figure 7: The comparison of eigenvalues(a) and eigenstates(b) obtained from the effective Hamiltonian and the quasiperiodic mosaic model. The black points, blue points and red points are the results from $\hat{H}$, $\hat{H}_{e}$ and $\hat{H}_{o}$ respectively. We choose $\lambda=10$ and $L=3\times 10^{3}$. $n$ labels the $n_{th}$ eigenstate.
@@ Appendix D Transformation between different basis
Consider the wave packet only propagating in the odd sites in the effective Hamiltonian, we can reduce above expressions: (1) for even sites:

### v2

@@ III Effective Hamiltonian in the large $\lambda$ limit
In the large quasiperiodic potential strength region, we have unveiled that the dynamics in the quasiperiodic mosaic lattice exhibits some special properties, including differences in density distribution between even and odd sites as well as time scales linearly depend on the strength of quasiperiodic potential. To understand these special dynamical behaviors, we shall provide an explanation by using an effective Hamiltonian in the subsequent discussion. Suppose that the quasiperiodic potential is sufficiently large, surpassing the strength of the hopping terms. We can apply the Schrieffer-Wolf (S-W) transformation method to construct the effective Hamiltonian for the quasiperiodic mosaic model. The effective Hamiltonian can provide a clear physical insight for understanding the special dynamical behaviors in the large quasiperiodic potential strength region.
@@ III.1 Derivation of the effective Hamiltonian
$\hat{H}=\textbf{c}^{\dagger}H\textbf{c},$ (14)
@@
$S_{mn}=$ $\sum_{p}[\frac{J}{\lambda_{p}}(\delta_{m,p}\delta_{n,p+1}+\delta_{m,p}\delta_{n,p-1})$ (15)
$-\frac{J}{\lambda_{p}}(\delta_{m-1,p}\delta_{n,p}+\delta_{m+1,p}\delta_{n,p})],$ (16)
@@
The subscript of $e$ or $o$ indicates the Hamiltonian defined on the even or odd sites. The details about S-W transformation can be found in Appendix B. To the first order, $\hat{H}_{e}$ and $\hat{H}_{o}$ are decoupled and their dynamics seem to be independent of each other. Next, we will study the properties the effective Hamiltonian and show how to use it to understand the special dynamical behaviors of quasiperiodic mosaic model.
@@ III.2 Properties of the effective Hamiltonian
First, we present the eigenstates and eigenvalues of the effective Hamiltonian, comparing them with the results obtained from exact diagonalization (ED). The results are illustrated in Fig. 5. In Fig. 5 (a), we depict the eigenvalues from the effective Hamiltonian and the quasiperiodic mosaic model, exhibiting good agreement between them. It is evident that the eigenvalues of $\hat{H}_{o}$ contribute to the energy levels of the band center around $E=0$, while the eigenvalues of $\hat{H}_{e}$ contribute to the energy levels apart from the band center. In Fig. 5 (b), we show the IPR for the effective Hamiltonian and quasiperiodic mosaic model. The effective Hamiltonian can capture the extended-localized properties of a large number of states. Although some eigenstates at the edge of the energy spectrum $\hat{H}_{o}$ deviate from the true eigenstates, we will show these eigenstates are localized and thus do not contribute to the dynamical behaviors.
Figure 5: The comparison of eigenvalues (a) and eigenstates (b) obtained from the effective Hamiltonian and the quasiperiodic mosaic model. The black points, blue points and red points are the results from $\hat{H}$, $\hat{H}_{o}$ and $\hat{H}_{e}$ respectively. We choose $\lambda=10$ and $L=377$. $n$ labels the $n$-th eigenstate.
@@
Figure 6: (a) IPR for eigenstates of $\hat{H}_{o}$ with different system sizes. We choose $\lambda=10$ and $L=377,~{}610,~{}987$. (b) Finite-size scaling analysis for different eigenstates.
(ii) $\hat{H}_{o}$ distributes on the odd sites. Both the onsite potential and hopping term have been modulated by the quasiperiodic potential equally. As shown in Fig. 6(a): the IPRs for the eigenstates of $\hat{H}_{o}$ in the middle of the energy spectrum is much smaller than that at the two sides of the spectrum, indicating that eigenstates in the middle and sides are extended and localized, respectively. To see it more clearly, in Fig. 6(b), we perform finite-size scaling analysis on typical eigenstates in different regions using the moments $D_{2}$ (Eq.5) to characterize the distribution information of the eigenstate. For our analysis, we take the logarithm of both sides of Eq.(5) and obtain
$-\ln(I_{2})/\ln(L)=-c/\ln(L)+D_{2}.$ (19)
We determine $D_{2}$ by conducting a linear fit to the curve in the two-dimensional space spanned by $-\frac{\ln(I_{2})}{\ln(L)}$ and $-\frac{1}{\ln(L)}$. We find that the fractal dimension of the typical eigenstate (red diamond markers) in the middle of the spectrum tends to 1 as $L\to\infty$, indicating that the eigenstate is extended. In contrast, the fractal dimension tends to 0 for the eigenstate (blue circle markers) at the edges of the spectrum, signifying that the eigenstate is localized. These results suggest the existence of mobility edges in the spectrum of $\hat{H}_{o}$.
From the effective Hamiltonian, we gain additional insights: (i) All the extended eigenstates can be traced back to the term $\hat{H}_{o}$. Despite the fact that the energy width of the extended states narrows as $\lambda$ increases, the proportion of extended states remains constant with the quasiperiodic potential strength. (ii) Two types of localized eigenstates exist in the quasiperiodic mosaic model. They distribute on different sites and have different physical origins. The first type of localized eigenstates originates from $\hat{H}_{e}$. These states are distributed at sites where the onsite potential is nonzero. Considering an eigenstate $|\psi_{j}\rangle$ with a localized center at the $j$-th site, its eigenvalue can be approximated as $\lambda_{j}+\frac{2J^{2}}{\lambda_{j}}$ which is proportional to quasiperiodic potential strength $\lambda$ in the large quasiperiodic potential strength case. The second type of localized eigenstates originates from $\hat{H}_{o}$. These eigenstates mainly distribute on sites where the potential is zero. The eigenvalues of this eigenstate are inversely proportional to the quasiperiodic potential strength $\lambda$.
@@
III.3 Understanding the wave packet dynamics from the effective Hamiltonian
@@
$e^{-i\hat{H}t}|\psi(0)\rangle$ $\approx e^{-i(\hat{H}_{e}+\hat{H}_{o})t}|\psi(0)\rangle$ (21)
@@
Since the wave packet dynamics is dominated by $\hat{H}_{o}$, the wave function only propagates along the odd sites in the dressed basis ($d^{\dagger}_{j}|0\rangle$). For the initial state with the Gaussian wave packet located at the center of the lattice, the particle number distribution $n^{d}_{j}=\langle\psi(t)|\hat{d}^{\dagger}_{j}\hat{d}_{j}|\psi(t)\rangle$ out of the initial wave packet only distributes on the odd sites. Nevertheless, according to the S-W transformation Eq.(17), the creation (annihilation) operator on the even sites in the dressed basis can be related to the operators on both even and odd sites in the original basis. Up to the second order, the density distribution on the even sites in the original basis can be explicitly expressed as
@@
$\frac{J^{2}}{\lambda_{j}^{2}}(\langle\psi(t)|\hat{d}^{\dagger}_{j-1}\hat{d}_{j+1}|\psi(t)\rangle+\langle\psi(t)|\hat{d}^{\dagger}_{j+1}\hat{d}_{j-1}|\psi(t)\rangle),$ (22)
@@
$-\frac{J^{2}}{2\lambda_{j+1}^{2}}(\langle\psi(t)|\hat{d}^{\dagger}_{j+2}\hat{d}_{j}|\psi(t)\rangle+\langle\psi(t)|\hat{d}^{\dagger}_{j}\hat{d}_{j+2}|\psi(t)\rangle).$ (23)
@@
From the expression of $\hat{H}_{o}$, we observe that $\frac{1}{\lambda}$ is an overall factor of the Hamiltonian of $\hat{H}_{o}$. It follows that the dynamics governed by $\hat{H}_{o}$ is invariant after a time scale $t\rightarrow t/\lambda$. This is consistent with the numerical results as shown in the inset of Fig.2. The wave packet dynamics governed by the effective Hamiltonian is also shown in Fig.1(e) with the cyan and magenta lines, exhibiting excellent agreement with the wave packet dynamics of the original Hamiltonian. Our results suggest that the effective Hamiltonian can describe the dynamical behaviors well for the mosaic model in the large quasiperiodic potential strength region.
IV Summary and discussion
In summary, we investigated the dynamics of a quasiperiodic mosaic model and unravel its peculiar dynamical behaviors in the large quasiperiodic potential region, exhibiting obvious different density distribution on even and odd sites. Our results also suggest that time scales of dynamical evolution are determined by the strength of the quasiperiodic potential linearly. Applying the S-W transformation, we derived an effective Hamiltonian in the limit of large quasiperiodic potential strength. The effective Hamiltonians can be divided into two decoupled Hamiltonians $\hat{H}_{e}$ and $\hat{H}_{o}$ defined on the even and odd sites, respectively. While all eigenstates of $H_{e}$ are localized, the eigenstates of $\hat{H}_{o}$ include both localized and extended eigenstates separated by mobility edge. The effective model describes the dynamical behaviors well for the mosaic model in the large quasiperiodic potential region. Our paper provides a clear picture for understanding the peculiar dynamical behaviors in the quasiperiodic mosaic lattice.
Although in this paper we only study the particular case of quasiperiodic mosaic models with $\kappa=2$ MosaicAA , our scheme can directly apply to study dynamics in other cases with $\kappa>2$. Considering the case of $\kappa=3$, we can still apply S-W transformation to derive an effective Hamiltonian in the limit of large quasiperiodic potential strength, which is described by two weakly coupled chains (see Appendix.E). We anticipate that the wavepacket dynamics will also exhibit distinct density distributions on the modulated and unmodulated sites.
@@ Appendix A Two-point correlation function
the correlation function can be written as
@@
Next we demonstrate how to calculate $\langle\hat{c}^{\dagger}_{a}\hat{c}_{b}\rangle$: the initial state $|\psi(0)\rangle$ is chosen as:
@@
To calculate the correlation function, we define two matrices $P^{A}$ and $P^{B}$. $P^{A}$ and $P^{B}$ are both $L\times(L/2+1)$ matrices and they satisfy:
@@ Appendix B S-W transformation
Then we can write the Hamiltonian as
@@ Appendix C Effectiveness of unitary transformation
In this paper, we investigate the impact of increasing the system size, denoted as $L$, on the effectiveness of the effective Hamiltonian under the same $\lambda$. Specifically, we choose $L=3\times 10^{3}$ and compare the eigenvalues and IPR obtained from the effective Hamiltonian and exact ED. The main results are depicted in Fig.7.
@@
Figure 7: The comparison of eigenvalues(a) and eigenstates(b) obtained from the effective Hamiltonian and the quasiperiodic mosaic model. The black points, blue points and red points are the results from $\hat{H}$, $\hat{H}_{e}$ and $\hat{H}_{o}$ respectively. We choose $\lambda=10$ and $L=3\times 10^{3}$. $n$ labels the $n$th eigenstate.
@@ Appendix D Transformation between different basis
Considering the wave packet only propagating in the odd sites in the effective Hamiltonian, we can reduce above expressions: (1) for even sites:
@@
Appendix E S-W transformation for $k=3$
Here, we show that the S-W transformation can also be applied in the case of $\kappa=3$. We choose the system size to be an integer multiple of $3$, and define a unit cell consisting of 3 sites, with the quasiperiodic potential placed on the second site. The Hamiltonian for the system is given by:
$\hat{H}=J\sum_{j}(\hat{c}_{j}^{\dagger}\hat{c}_{j+1}+\text{H.c.})+2\sum_{j}\lambda_{j}\hat{n}_{j},$
where $\lambda_{j}$ is defined as:
$\lambda_{j}=\begin{cases}\lambda\cos(2\pi\omega j+\phi),&\text{if }\mod(j+1,\kappa)=0,\\ 0,&\text{otherwise}.\end{cases}$
with $\kappa=3$. Following a similar procedure, we introduce an anti-Hermitian matrix:
$S_{mn}=$ $\sum_{p}\frac{J}{\lambda_{p}}\left(\delta_{m,p}\delta_{n,p+1}+\delta_{m,p}\delta_{n,p-1}\right)$
$-\frac{J}{\lambda_{p}}\left(\delta_{m-1,p}\delta_{n,p}+\delta_{m+1,p}\delta_{n,p}\right),$
where $\mod(p+1,3)=0$. This transformation allows us to decompose the system into two weakly coupled chains, with the effective Hamiltonian expressed as:
$\hat{H}=\hat{H}_{E}+\hat{H}_{L}+\hat{H}_{C},$
where the components are:
[…]
$\hat{H}_{L}=\sum_{j=2,5,8,\dots}^{L-1}\left(\lambda_{j}+\frac{2J^{2}}{\lambda_{j}}\right)\hat{d}_{j}^{\dagger}\hat{d}_{j},$
[…]
The Hamiltonian $\hat{H}_{E}$ includes weak disorder potentials and hopping terms. This part contributes to extended eigenstates, although the dynamics are more complex because the quasiperiodic potential strength is not a global factor. The term $\hat{H}_{L}$ contributes to localized eigenstates, while $\hat{H}_{C}$ represents weak coupling between the two chains. The overall dynamics of the system are determined by the interplay of these terms.